OpenStudy (anonymous):

Prove the identity: Cos^2x-sin^2x=1-2sin^2x

OpenStudy (zarkon):

\[\cos^2(x)-\sin^2(x)\] \[=\cos^2(x)+\sin^2(x)-\sin^2(x)-\sin^2(x)=\cdots\]